what is the calculation and errors of simple pendulum expriment ?


1
Expert's answer
2022-03-23T07:15:37-0400

The period of oscillation of a simple pendulum is given by

"T=2\pi\sqrt{\frac{l}{g}}"

In experiment we use a simple pendulum to calculate the acceleration due gravity

"g=\frac{4\pi^2l}{T^2}"

The uncertainty of acceleration is given by

"\frac{\Delta g}{g}=\sqrt{\left(\frac{\Delta l}{l}\right)^2+\left(\frac{2\Delta T}{T}\right)^2}"